OpenStudy (anonymous):

the bakery each day makes pastries 25% are choclate do

OpenStudy (anonymous):

sorry it was 25% are chocolate doughnuts 37.5 are butter cookies 25% are breakfast buns and the rest are the daily specail there are 90 pastries

OpenStudy (anonymous):

how many pastries are made each day?

OpenStudy (anonymous):

what percent of each day's pastreis are the daily specail

OpenStudy (anonymous):

and How many of the daily specail pastries are made

OpenStudy (anonymous):

please REPLY!!!!!!!!!!!!!!!!!! SOOOOOOOOOOOOOOOOOOOOOOOONNNNNNNNNN

OpenStudy (texaschic101):

doesn't your question specify that there are 90 pastries made each day ?

OpenStudy (anonymous):

well no

OpenStudy (anonymous):

there are 90 breakfast buns

OpenStudy (texaschic101):

no...there are 25% of 90 breakfast buns. There are 90 pastries. 25% are chocolate doughnuts 0.25(90) = 22.5 25% are breakfast buns = 22.5 37.5 are butter cookies 22.5 + 22.5 + 37.5 = 82.5 that leaves 90 - 82.5 = 7.5 daily specials 7.5/90 = .083 = 8.3% are daily specials

OpenStudy (anonymous):

thanx soooooooooooooooooo much

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):

oh also how many pastreis are made each day?

OpenStudy (texaschic101):

90 pastries are made each day

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

bye

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

ya